OpenStudy (anonymous):

Find an equation of the tangent curve at the given point? \(\ \large y=sin(sinx), (\pi,0) \).

OpenStudy (ash2326):

@Study23 Slope of the tangent at the point \((\pi, 0)\) to the curve \(y=\sin (\sin x)\) is given by \[\frac{dy}{dx}_{(\pi, 0)}\] Can you find ? \[\frac{dy}{dx}\]

OpenStudy (anonymous):

Could you help me with that? I'm not sure how to use the values to find the derivative, @ash2326

OpenStudy (ash2326):

Just find the derivative, we'll use the values later.

OpenStudy (anonymous):

Okay, one moment while I find the derivative

OpenStudy (ash2326):

ok

OpenStudy (anonymous):

\(\ \Huge \text{So, I got: } y'=cos^2x \)

OpenStudy (anonymous):

\(\ \large \text{How do I proceed from here?} \)

OpenStudy (ash2326):

\[\frac{dy}{dx}=\frac{d}{dx} (\sin (\sin x))\] \[\frac{dy}{dx}=\cos (\sin x)\times\frac{d}{dx} \sin (x)\]

OpenStudy (anonymous):

? What I did: \(\ \large y=sin(sinx) \) \(\ \large y'=cos\frac{d}{dx}(sinx)\) \(\ \large y' = cos \times cosx,\) \(\ \large \text{So, } cos^2x \)

OpenStudy (ash2326):

you'd get \[\cos (\sin x)\times \cos x\]

OpenStudy (anonymous):

Okay, I got that this time... So now what?

OpenStudy (ash2326):

now put x=\(\pi\) here to find the slope of the tangent

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

So, -1*-1=1? That's the slope?